OpenStudy (biohazard9064):

Please Help. Solve -4(x + 1) - 3 = -3(x - 4).

OpenStudy (biohazard9064):

A. -19 B. 19 C. 2 D. 0

OpenStudy (midhun.madhu1987):

-4(x + 1) - 3 = -3(x - 4) Open the bracket : -4x - 4 - 3 = -3x + 12 now move all x to left side and all numbers to right side and simplify

OpenStudy (biohazard9064):

@midhun.madhu1987 how do I do that?

OpenStudy (biohazard9064):

so -4-4-3=3x+12 how would I simplify ? by using -4?

OpenStudy (midhun.madhu1987):

-4x + 3x = 12 + 4 + 3 Now simplify this..

OpenStudy (biohazard9064):

-1x=19

OpenStudy (biohazard9064):

@midhun.madhu1987 would it be 19

OpenStudy (midhun.madhu1987):

x = -19

OpenStudy (biohazard9064):

thanks